OpenStudy (anonymous):

HELPPPP!!! Solve 5^(x + 5) = 9^x

OpenStudy (mathstudent55):

Take log of both sides

OpenStudy (anteater):

Start by taking the log or natural log of both sides

OpenStudy (anonymous):

i put this in on my calculator log 9 ÷ log 5

OpenStudy (anteater):

Ok, so log 9 divided by log 5 equals what?

OpenStudy (anonymous):

1.365

OpenStudy (anteater):

Ok, and then what does the other side of your equation look like?

OpenStudy (anonymous):

should it look like this ? x+5 = 1.365^x

OpenStudy (anteater):

Yes, good :D

OpenStudy (anteater):

Oops

OpenStudy (anteater):

1.365 times x on the right

OpenStudy (anonymous):

um can you type it out please ? (: i dont understand you

OpenStudy (anteater):

So x + 5 = 1.365x

OpenStudy (anonymous):

ohhhh my bad lol,i see now

OpenStudy (anonymous):

then you subtract 5 from the right side right ?

OpenStudy (anteater):

Yes, from left and right sides :)

OpenStudy (anteater):

So x = 1.365x - 5

OpenStudy (anonymous):

alrighty, then youuuuuu divide x ?

OpenStudy (anteater):

I think you might want to subtract 1.365x from both sides, to get all the x's on one side of the equation.

OpenStudy (mathstudent55):

This is how you should start it: